Bluegold, Inc. reported total shareholders’

equity at $40,000 and net income was

$30,000. The company currently has 10,000

shares outstanding. The price per share for

Bluegold is currently $25. What is the firm’s

P/E ratio? What is the firm’s P/B ratio? What

do these numbers indicate about how

shareholders view Bluegold?
Answer #1

Earnings Per Share, EPS = Net Income/Number of shares outstanding

EPS = 30,000/10,000 = $3

P/E = Price/EPS

P/E = 25/3 = 8.3333333333

PE ratio indicates how much each share of the company costs per dollar of the net income.

Book value of equity per share = Equity/Number of shares outstanding

Book value = 40,000/10,000 = $4

P/B = 25/4 = 6.25

This indicates that if the firm were to liquidate its assets, then each shareholder will get $6.25 per $1 of shareholders' equity.

Can you please upvote? Thank You :-)